# Grounded GUI Understanding for Vision Based Spatial Intelligent Agent: Exemplified by Virtual Reality Apps

### v3

@@ 3.1. Problem Formulation of IGE Detection
Before jumping into the details of our approach, we first introduce how we formulate the IGE detection problem. The basic IGE detection unit is an individual VR app scenario, which lies in the user’s field of view, at a specific time. Unlike mobile application GUIs or web app GUIs, which are entirely 2D, HMD-based VR apps typically render two 2D GUIs for both eyes of users, creating an illusion of depth and making users feel stereoscopic 3D (S3D) sense (Penn and Hout, 2018; Elmqaddem, 2019). These scenes, presented to each eye, are projections of the 3D virtual world onto two 2D ”picture planes” (user’s eyes), mathematically represented by the function: $P_{3D\rightarrow 2D}:UI_{3D}\rightarrow(UI_{left},UI_{right}),$ where $UI_{3D}$ represents the 3D scene in the VR environment, and $UI_{left}$ and $UI_{right}$ represent the corresponding 2D projections rendered for the left and right eyes, respectively. Each one-eye scene can be recovered using the other eye’s scene (Li et al., 2024). Therefore, the S3D IGEs can be detected by analyzing the 2D projection from a single eye. This allows us to simplify the VR IGE detection problem to detecting IGEs in the 2D scene of any individual eye.
To facilitate downstream SE tasks such as boosting automated GUI testing, each IGE is identified not only by its interactability status but also by its location and associated semantic labels, enabling a more fine-grained understanding of the IGEs. Let $GE$ represent the set of GUI elements in the VR scene under analysis, i.e., $UI$. Given that the 3D scene $UI_{3D}$ is projected into 2D for each eye, the right-eye scene can be denoted as: $P_{3D\rightarrow 2D}(UI_{right}):UI_{3D}\rightarrow UI_{right}.$ The detection function $D:GE_{2D}\rightarrow\{Non-interactable,Interactable\}$ maps each GUI element to a binary interactability determination. The function $D$ operates on the 2D projections of the 3D GUI elements, denoted as:
@@
Here, $C$ denotes the set of semantic contexts that facilitate context-sensitive analysis. We use a bounding box $B$ to specify the spatial location of each IGE, which is the smallest unrotated rectangle containing that IGE and can be described by its upper-left corner coordinates, width, and height. Let $S$ denote the set of semantic labels associated with each GUI element, where the labels and their granularity depend on the specific VR scene. As presented above, the problem can be simplified to detection on the 2D scene of any individual eye. If we use the right-eye scene for demonstration, the IGE detection problem can be formulated as:
@@ 3.2. Overview of Orienter
Figure 4 illustrates our cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps, short as Orienter. To address the three challenges (section 1), Orienter employs a feedback-driven approach that mirrors human behavior, prioritizing semantic comprehension before initiating detection; Orienter also enables iterative refinement and validation through a feedback loop. Such designs embody the principle of “looking before leaping”.
Orienter consists of three primary components: (1) Semantic Context Comprehension: To enable unsupervised, context-sensitive IGE detection, Orienter first synthesizes both the global context (e.g., app genre, storyline, interaction mechanisms) and the local context (i.e., the user’s current VR field of view). Leveraging the LMM, Orienter performs in-context reasoning to establish a comprehensive understanding of the VR scene, forming the basis for subsequent detection. (2) Reflection-Guided IGE Candidate Detection: Orienter then mines multi-perspective characteristics of GUI elements, and uses these characteristic descriptions for detecting and locating potential IGEs. The framework iteratively refines these candidates through a feedback loop, performing reflection-driven validation to reduce hallucination. (3) Context-Sensitive Interactability Classification: Finally, Orienter predicts the interactability of the detected IGE candidates by incorporating the semantic context from the first stage. Using chain-of-thought reasoning within the LMM, Orienter classifies interactability without requiring labeled training data, leveraging a knowledge-transfer paradigm from pretrained models in other domains.
@@ 3.3. Module I: Semantic Context Comprehension
The interactability of GUI elements in VR apps is highly dependent on the semantic context of the VR app content and scenarios. To tackle this challenge, Orienter understands and analyzes the target VR app’s context beforehand to enable context-sensitive detection. The semantic context is analyzed comprehensively, from (1) both global (overall app context) and local (current VR scenario) granularities, and (2) both natural language (texts) to vision (images) modalities.
@@
The global context captures the overall semantic information of the VR app under analysis. To find out which global context attributes have influences on IGE interactability, two authors randomly sample 30 real-world VR applications from SteamVR and perform manual inspection following the open coding procedure (Creswell and Poth, 2016). At last, we reach a consensus and locate the following key attributes: app names, genres, content themes, VR device supports, ways of gameplay, possible interaction mechanisms, and language information. Such context can be extracted from the detailed information page on official VR app stores, but effective extraction can be troublesome due to the free structure natural language format of app details. Hence, Orienter leverages LMM to perform the extraction after the app details are crawled automatically. Prompt PI.1 in Figure 2 shows the prompt template.
@@
Local context represents the semantic context within the current VR scenario, i.e., all VR content within the user’s current field of view. Under the extracted global context of app content from section 3.3.1, Orienter further guides LMM to perform in-context visual question answering, to digest and summarize the rendered VR scene as local context. The guidance of global context ensures that the interpretation of the VR scene is informed by the narrative and purpose of the VR app under analysis. Specifically, the LMM processes the composite input (screenshot and global context) and outputs a summary of all GUI elements and the background within the scene. Prompt PI.2 in Figure 2 shows the prompt template.
@@ 3.4. Module II: Reflection-Directed IGE Candidate Detection
This module aims at recognizing and localizing IGE candidates within the VR scene, based on the captured semantic contexts from Module I. LMM can reveal more semantic hints than traditional methods, making it a better backbone for our approach. However, LMM tends to generate semantically-incorrect IGEs with inaccurate locations. This is because VR IGE detection requires precise spatial perception and visual-semantic alignment capabilities, and VR IGE categories are open-vocabulary and heterogeneous (as presented in the challenges and motivational examples). Orienter conquer these challenges through a reflection-directed IGE candidate detection loop:
@@
iii). Inspired by the human cognitive process, when trying to find and locate IGE candidates, using a semantic label only will make the detection process unrobust, e.g., may find the wrong element or find the wrong locations. Adding more descriptions about the target element (like color, size, shape, relevant locations, etc.) makes the detection more effective and robust. It is similar for To boost the effectiveness and robustness of i) and ii), Orienter extends a direct LMM detector to a multi-perspective characteristics miner (section 3.4.1).
@@ 3.4.1. Multi-Perspective Characteristics Mining
In this module, Orienter firstly identifies all IGE candidates in the VR scene, and then mines their diverse characteristics from different perspectives (e.g., size, color, shape, relevant location with other objects, etc.). Then these characteristics will form a characteristics description, which can uniquely identify the corresponding IGE candidate in the present figure in the described detection module. Figure 7 shows the prompt templates and examples.
Figure 7. Prompt templates and examples for Module II.①2
II.①(a): IGE Candidate Recognition (Prompt PII.1). Orienter engages in the recognition of IGE candidates within the VR scene. With the semantic context from Module I (section 3.3), Orienter analyzes the visual semantic cues presented in the environment, finding out all IGE candidates.
II.①(b): Characteristics Dimension Recognition (Prompt PII.2). Following the identification of IGE candidates, Orienter embarks on recognizing the characteristics dimensions of them, which makes them distinguishable from other GUI elements. This involves dissecting the unique characteristics attributes and properties related to the visual appearance (e.g., color, size, shape, etc.), functionalities, interaction mechanisms, and relative locations of each GUI element.
@@
II.①(d) In-Context Characteristics Reasoning (Prompt PII.4). Then Orienter conducts characteristics reasoning, querying the LMM to analyze the VR scene and find out the answers for formulated characteristic questions. This step integrates the local and global contexts previously established, allowing the framework to interpret the responses within the broader narrative and functional scope of the VR application. Through this in-context reasoning, Orienter synthesizes a comprehensive understanding of each IGE candidate’s characteristics. Orienter then generates characteristics descriptions (CDs), which describe the GUI elements based on these characteristics. Figure 7 shows an example of a characteristics description, “A white and black small spherical baseball with grey shadows on the bottom half and glowing trail across the center of the scene”.
@@ 3.4.2. Described IGE Candidate Detection
Figure 8. The framework of described IGE candidate detection module (VFMD)
@@
Orienter detects and localizes IGEs using the descriptions of the crafted characteristics generated in the previous step, inspired by Shen’s work (Shen et al., 2024). Orienter employs a language-model-aligned visual foundation model (VFMD). Figure 9 shows the framework structrure. VFMD consists of a vision backbone for extracting visual features from VR scenes and a language model for generating text embeddings from CDs. Text embedding s are aggregated into sentence-level representations and are fused with visual embeddings using a cross-modality transformer encoder. Object queries conditioned on the text are processed by a transformer decoder, producing object embeddings. A final visual-language alignment module predicts the correct pairings of IGE candidate locations and CDs. The backbone language model, trained on a vast corpus of data, can generalize to the diverse and unlimited categories of GUI elements present in VR scenarios. The localization process cross-references (or aligns with) the textual CDs with the visual input from the VR scene. By doing so, VFMD successfully locates IGE candidates. Here is an example to demonstrate the output of this module, as shown in Figure 9.
@@ 3.4.3. Feedback-Directed Reflection
Mirroring-Based Regional Verification for Successfully Detected CDs (Prompt PII.5). Conversely, for successfully detected IGEs, we design a method of regional zooming and mirroring. This involves cropping the detected GUI element within its bounding box into a new figure, and Orienter lets LMM conduct a comparative analysis against the original VR scene. Such an approach facilitates a granular verification of the detection’s accuracy.
@@ 3.5. Module III: Context-Sensitive Interactability Classification
After establishing the presence and location of GUI elements in the VR environment through the first two modules, Orienter proceeds to analyze their interactivity. This module determines which GUI elements in the VR scene can be interacted with. Such information is useful for many downstream SE tasks, including boosting automated testing performance (detailed in section 6.3).
Figure 11 shows the prompt template PIII. To classify interactability, Orienter guides LMM towards a chain-of-thought in-context reasoning process, within the semantic context captured in Module I. This process mimics a VR player’s thought pattern in distinguishing between GUI elements that are mere visual elements and those that afford interaction. For instance, a tree in a tree-planting game (Figure 1(c)) is interactable, serving as a GUI element to be manipulated by the player, whereas the same tree would be non-interactable background scenery in a fishing game (Figure 1(d)). Orienter also provides demonstration examples in the prompt, illustrating the reasoning required for differentiating between interactable and non-interactable GUI elements.
@@ 4. Dataset Construction
To the best of our knowledge, there is no existing GUI dataset for IGE detection in VR apps. Therefore, we build a dataset to verify the usefulness and effectiveness of Orienter.
We recruit a team of 13 annotators, spending more than three months in the collection and annotation of GUI images from various VR apps. These annotators have a minimum of two years of computer science or electrical engineering background, and mostly with experience in video games. The annotators first interact with GUI elements in various VR apps while their views are recorded. They then select GUI images from these recordings and annotate IGEs. To enhance the quality of the dataset, we train the annotators in the use of VR equipment and the labeling of IGEs, and we ask them to follow several guidelines during data collection and annotation: (1) attempt to identify all IGEs using every possible interaction method in each VR scene, and (2) categorize the semantics of IGEs with appropriate granularity based on their context.
4.1. Collection of VR Apps
We first collect VR apps from the Steam app store (web, 2023c), a comprehensive repository with a wide variety of VR content. This collection yielded a total of 4,610 VR-Only apps, which includes software (official type label on Steam referring to non-games) and games. To obtain a statistically significant sample size with a 95% confidence level and a 10% margin of error, we randomly sample 102 apps for analysis to cover a broader spectrum of categories, contexts, and interaction paradigms, covering 245 community-generated genres on Steam, demonstrating their diversity and representativeness.
4.2. Collection of GUI Images from VR Apps
In this step, the annotators engage with the selected VR apps, interacting with all GUI components within the VR scene in every possible way, as described in section 2.1.2, while recording their stereo view. The annotators then select and save GUI images that encompass all VR scenes and GUI elements they explored from the recordings. We crop the right-eye images from the stereo-view recordings, as only one side of the view is needed, as explained in section 3.1.
@@ 4.2. Collection of GUI Images from VR Apps
4.3. Annotation of IGEs in GUI Images
In this step, the annotators identify IGEs in GUI images and label their locations with bounding boxes as well as their semantics within the context. As discussed in section 1, the interactability and semantics of GUI elements heavily depend on the diverse contexts of VR apps, making it challenging to categorize all GUI elements with a finite set of predefined categories.
To address this challenge, we ask the annotators to create categories. Specifically, they are instructed to classify IGEs with appropriate granularity based on their experience, to better describe the semantics of IGEs in concrete contexts. For example, in the VR game, VR The Diner Duo (web, 2023d), fish is only one kind of ingredient to make burgers, as shown in 12(a). Different individuals of fish are semantically the same in this game. Therefore, a coarse-grained fish category is adequate to describe it. However, in the game Munch VR (web, 2023b), players need to control a fish to eat other smaller fish while avoiding hunted by larger fish, as shown in 12(b). In this scenario, fish individuals diverge semantically according to size and appearance, necessitating the introduction of more fine-grained categories such as large fish and small fish to capture their semantic distinctions.

### v4

@@ 3.1. Problem Formulation of IGE Detection
Before jumping into the details of our approach, we first introduce how we formulate the IGE detection problem. The basic IGE detection unit is an individual XR app scenario, which lies in the user’s field of view, at a specific time. Unlike mobile application GUIs or web app GUIs, which are entirely 2D, HMD-based XR apps typically render two 2D GUIs for both eyes of users, creating an illusion of depth and making users feel stereoscopic 3D (S3D) sense (Penn and Hout, 2018; Elmqaddem, 2019). These scenes, presented to each eye, are projections of the 3D virtual world onto two 2D ”picture planes” (user’s eyes), mathematically represented by the function: $P_{3D\rightarrow 2D}:UI_{3D}\rightarrow(UI_{left},UI_{right}),$ where $UI_{3D}$ represents the 3D scene in the XR environment, and $UI_{left}$ and $UI_{right}$ represent the corresponding 2D projections rendered for the left and right eyes, respectively. Each one-eye scene can be recovered using the other eye’s scene (Li et al., 2024). Therefore, the S3D IGEs can be detected by analyzing the 2D projection from a single eye. This allows us to simplify the XR IGE detection problem to detecting IGEs in the 2D scene of any individual eye.
To facilitate downstream SE tasks such as boosting automated GUI testing, each IGE is identified not only by its interactability status but also by its location and associated semantic labels, enabling a more fine-grained understanding of the IGEs. Let $GE$ represent the set of GUI elements in the XR scene under analysis, i.e., $UI$. Given that the 3D scene $UI_{3D}$ is projected into 2D for each eye, the right-eye scene can be denoted as: $P_{3D\rightarrow 2D}(UI_{right}):UI_{3D}\rightarrow UI_{right}.$ The detection function $D:GE_{2D}\rightarrow\{Non-interactable,Interactable\}$ maps each GUI element to a binary interactability determination. The function $D$ operates on the 2D projections of the 3D GUI elements, denoted as:
@@
Here, $C$ denotes the set of semantic contexts that facilitate context-sensitive analysis. We use a bounding box $B$ to specify the spatial location of each IGE, which is the smallest unrotated rectangle containing that IGE and can be described by its upper-left corner coordinates, width, and height. Let $S$ denote the set of semantic labels associated with each GUI element, where the labels and their granularity depend on the specific XR scene. As presented above, the problem can be simplified to detection on the 2D scene of any individual eye. If we use the right-eye scene for demonstration, the IGE detection problem can be formulated as:
@@ 3.2. Overview of Orienter
Figure 4 illustrates our cOntext-sensitive inteRactable GUI ElemeNT detection framework for Extended Reality apps, short as Orienter. To address the three challenges (section 1), Orienter employs a feedback-driven approach that mirrors human behavior, prioritizing semantic comprehension before initiating detection; Orienter also enables iterative refinement and validation through a feedback loop. Such designs embody the principle of “looking before leaping”.
Orienter consists of three primary components: (1) Semantic Context Comprehension: To enable unsupervised, context-sensitive IGE detection, Orienter first synthesizes both the global context (e.g., app genre, storyline, interaction mechanisms) and the local context (i.e., the user’s current XR field of view). Leveraging the LMM, Orienter performs in-context reasoning to establish a comprehensive understanding of the XR scene, forming the basis for subsequent detection. (2) Reflection-Guided IGE Candidate Detection: Orienter then mines multi-perspective characteristics of GUI elements, and uses these characteristic descriptions for detecting and locating potential IGEs. The framework iteratively refines these candidates through a feedback loop, performing reflection-driven validation to reduce hallucination. (3) Context-Sensitive Interactability Classification: Finally, Orienter predicts the interactability of the detected IGE candidates by incorporating the semantic context from the first stage. Using chain-of-thought reasoning within the LMM, Orienter classifies interactability without requiring labeled training data, leveraging a knowledge-transfer paradigm from pretrained models in other domains.
@@ 3.3. Module I: Semantic Context Comprehension
The interactability of GUI elements in XR apps is highly dependent on the semantic context of the XR app content and scenarios. To tackle this challenge, Orienter understands and analyzes the target XR app’s context beforehand to enable context-sensitive detection. The semantic context is analyzed comprehensively, from (1) both global (overall app context) and local (current XR scenario) granularities, and (2) both natural language (texts) to vision (images) modalities.
@@
The global context captures the overall semantic information of the XR app under analysis. To find out which global context attributes have influences on IGE interactability, two authors randomly sample 30 real-world VR applications from SteamVR and perform manual inspection following the open coding procedure (Creswell and Poth, 2016). At last, we reach a consensus and locate the following key attributes: app names, genres, content themes, VR device supports, ways of gameplay, possible interaction mechanisms, and language information. Such context can be extracted from the detailed information page on official XR app stores, but effective extraction can be troublesome due to the free structure natural language format of app details. Hence, Orienter leverages LMM to perform the extraction after the app details are crawled automatically. Prompt PI.1 in Figure 2 shows the prompt template.
@@
Local context represents the semantic context within the current XR scenario, i.e., all XR content within the user’s current field of view. Under the extracted global context of app content from section 3.3.1, Orienter further guides LMM to perform in-context visual question answering, to digest and summarize the rendered XR scene as local context. The guidance of global context ensures that the interpretation of the XR scene is informed by the narrative and purpose of the XR app under analysis. Specifically, the LMM processes the composite input (screenshot and global context) and outputs a summary of all GUI elements and the background within the scene. Prompt PI.2 in Figure 2 shows the prompt template.
@@ 3.4. Module II: Reflection-Directed IGE Candidate Detection
This module aims at recognizing and localizing IGE candidates within the XR scene, based on the captured semantic contexts from Module I. LMM can reveal more semantic hints than traditional methods, making it a better backbone for our approach. However, LMM tends to generate semantically-incorrect IGEs with inaccurate locations. This is because XR IGE detection requires precise spatial perception and visual-semantic alignment capabilities, and XR IGE categories are open-vocabulary and heterogeneous (as presented in the challenges and motivational examples). Orienter conquer these challenges through a reflection-directed IGE candidate detection loop:
@@
iii). Inspired by the human cognitive process, when trying to find and locate IGE candidates, using a semantic label only will make the detection process unrobust, e.g., may find the wrong element or find the wrong locations. Adding more descriptions about the target element (like color, size, shape, relevant locations, etc.) makes the detection more effective and robust. To boost the effectiveness and robustness of i) and ii), Orienter extends a direct LMM detector to a multi-perspective characteristics miner (section 3.4.1).
@@ 3.4.1. Multi-Perspective Characteristics Mining
In this module, Orienter firstly identifies all IGE candidates in the XR scene, and then mines their diverse characteristics from different perspectives (e.g., size, color, shape, relevant location with other objects, etc.). Then these characteristics will form a characteristics description, which can uniquely identify the corresponding IGE candidate in the present figure in the described detection module. Figure 6 shows the prompt templates and examples.
Figure 6. Prompt templates and examples for Module II.①2
@@
II.①(a): IGE Candidate Recognition (Prompt PII.1). Orienter engages in the recognition of IGE candidates within the XR scene. With the semantic context from Module I (section 3.3), Orienter analyzes the visual semantic cues presented in the environment, finding out all IGE candidates.
II.①(b): Characteristics Dimension Recognition (Prompt PII.2). Following the identification of IGE candidates, Orienter embarks on recognizing the characteristics dimensions of them, which makes them distinguishable from other GUI elements. This involves dissecting the unique characteristics attributes and properties related to the visual appearance, functionalities, interaction mechanisms, and relative locations of each element.
@@
II.①(d) In-Context Characteristics Reasoning (Prompt PII.4). Then Orienter conducts characteristics reasoning, querying the LMM to analyze the XR scene and find out the answers for formulated characteristic questions. This step integrates the local and global contexts previously established, allowing the framework to interpret the responses within the broader narrative and functional scope of the XR application. Through this in-context reasoning, Orienter synthesizes a comprehensive understanding of each IGE candidate’s characteristics. Orienter then generates characteristics descriptions (CDs), which describe the GUI elements based on these characteristics. Figure 6 shows an example of a characteristics description, “A white and black small spherical baseball with grey shadows on the bottom half and glowing trail across the center of the scene”.
@@ 3.4.2. Described IGE Candidate Detection
Figure 8. The framework of the described IGE candidate detection module (VFMD)
Orienter detects and localizes IGEs using the descriptions of the crafted characteristics generated in the previous step, inspired by Shen’s work (Shen et al., 2024). Orienter employs a language-model-aligned visual foundation model (VFMD). Figure 8 shows the framework structrure.
@@
VFMD consists of a vision backbone for extracting visual features from XR scenes and a language model for generating text embeddings from CDs. Text embedding s are aggregated into sentence-level representations and are fused with visual embeddings using a cross-modality transformer encoder. Object queries conditioned on the text are processed by a transformer decoder, producing object embeddings. A final visual-language alignment module predicts the correct pairings of IGE candidate locations and CDs. The backbone language model, trained on a vast corpus of data, can generalize to the diverse and unlimited categories of GUI elements present in XR scenarios. The localization process cross-references (or aligns with) the textual CDs with the visual input from the XR scene. By doing, VFMD successfully locates IGE candidates. Figure 9 is an example showing the output of this module.
@@ 3.4.3. Feedback-Directed Reflection
Mirroring-Based Regional Verification for Successfully Detected CDs (Prompt PII.5). Conversely, for successfully detected IGEs, we design a method of regional zooming and mirroring. This involves cropping the detected GUI element within its bounding box into a new figure, and Orienter lets LMM conduct a comparative analysis against the original XR scene. Such an approach facilitates a granular verification of the detection’s accuracy.
@@ 3.5. Module III: Context-Sensitive Interactability Classification
After establishing the presence and location of GUI elements in the XR environment through the first two modules, Orienter proceeds to analyze their interactivity. This module determines which GUI elements in the XR scene can be interacted with. Such information is useful for many downstream SE tasks, including boosting automated testing performance (detailed in section 6.3).
Figure 11 shows the prompt template PIII. To classify interactability, Orienter guides LMM towards a chain-of-thought in-context reasoning process, within the semantic context captured in Module I. This process mimics a XR player’s thought pattern in distinguishing between GUI elements that are mere visual elements and those that afford interaction. For instance, a tree in a tree-planting game (Figure 1(c)) is interactable, serving as a GUI element to be manipulated by the player, whereas the same tree would be non-interactable background scenery in a fishing game (Figure 1(d)). Orienter also provides demonstration examples in the prompt, illustrating the reasoning required for differentiating between interactable and non-interactable GUI elements.
@@ 4. Dataset Construction
To the best of our knowledge, there is no existing GUI dataset for IGE detection in XR apps. Therefore, we build a dataset to verify the usefulness and effectiveness of Orienter.
We recruit a team of 13 annotators, spending more than three months in the collection and annotation of GUI images from various XR apps. These annotators have a minimum of two years of computer science or electrical engineering background, and mostly with experience in video games. The annotators first interact with GUI elements in various XR apps while their views are recorded. They then select GUI images from these recordings and annotate IGEs. To enhance the quality of the dataset, we train the annotators in the use of XR equipment and the labeling of IGEs, and we ask them to follow several guidelines during data collection and annotation: (1) attempt to identify all IGEs using every possible interaction method in each XR scene, and (2) categorize the semantics of IGEs with appropriate granularity based on their context.
4.1. Collection of XR Apps
We first collect XR apps from the Steam app store (web, 2023c), a comprehensive repository with a wide variety of XR content. This collection yielded a total of 4,610 VR-Only apps, which includes software (official type label on Steam referring to non-games) and games. To obtain a statistically significant sample size with a 95% confidence level and a 10% margin of error, we randomly sample 102 apps for analysis to cover a broader spectrum of categories, contexts, and interaction paradigms, covering 245 community-generated genres on Steam, demonstrating their diversity and representativeness.
4.2. Collection of GUI Images from XR Apps
In this step, the annotators engage with the selected XR apps, interacting with all GUI components within the XR scene in every possible way, as described in fig. 2, while recording their stereo view. The annotators then select and save GUI images that encompass all XR scenes and GUI elements they explored from the recordings. We crop the right-eye images from the stereo-view videos as only one side of the view is needed, as explained in section 3.1.
4.3. Annotation of IGEs in GUI Images
@@
The annotators then identify IGEs in GUI images and label their locations with bounding boxes as well as their semantics within the context. As discussed in section 1, the interactability and semantics of GUI elements heavily depend on the diverse contexts of XR apps, making it challenging to categorize all GUI elements with a finite set of predefined categories.
To address this challenge, we ask the annotators to create categories. Specifically, the annotators are instructed to classify IGEs with appropriate granularity based on their experience to better describe the semantics of IGEs in concrete contexts. For example, in the XR game, VR The Diner Duo (web, 2023d), fish is only one kind of ingredient to make burgers, as shown in 12(a). Different individuals of fish are semantically the same in this game. Therefore, a coarse-grained fish category is adequate to describe it. However, in the game Munch VR (web, 2023b), players need to control a fish to eat other smaller fish while avoiding hunted by larger fish, as shown in 12(b). In this scenario, fish individuals diverge semantically according to size and appearance, necessitating the introduction of more fine-grained categories such as large fish and small fish to capture their semantic distinctions.